# A proof of a conjecture of Erdős and Gyárfás on monochromatic path covers

### v1

@@ 2 Preliminaries
Let $C=160000$, $n_{0}=C^{10}$ and $n>n_{0}$. Define $\alpha_{n}=9C/\sqrt[4]{n}$, and suppose $\chi\in\mathcal{X}_{n}$ is such that $f(n,\chi)>\sqrt{n}$. By Proposition 3.4, we have $f(m)<\sqrt{m}+C$ for every $m\in\mathbb{N}$. Applying Lemma 3.2 with $C_{1}=C$ and $C_{2}=0$, we see that without loss of generality, there exists a blue path $P$ such that the size of $Y=[n]\setminus V(P)$ is at most $(1+\alpha_{n})\sqrt{n}$ and with the additional property that every vertex in $Y$ has less than $2C\sqrt{n}$ blue neighbours in $X=V(P)$. By Lemma 3.3, if $|Y_{0}|\leq(1-2\alpha_{n})\sqrt{n}$ or $|Y|\leq\sqrt{n}-1$, then $f(n,\chi)\leq\sqrt{n}$, a contradiction. Let us therefore assume that $|Y_{0}|>(1-2\alpha_{n})\sqrt{n}$ and $|Y|>\sqrt{n}-1$.
We define $H$ as the bipartite graph $H$ on vertex set $X\cup Y$ whose edges are the red edges between $X$ and $Y$ under $\chi$. Let $Y_{1}=Y\setminus Y_{0}$, $X_{0}=\{x\in X:d_{H}(x)=|Y|\}$, and $X_{1}=X\setminus X_{0}$. We have $|Y_{1}|=|Y|-|Y_{0}|<(1+\alpha_{n})\sqrt{n}-(1-2\alpha_{n})\sqrt{n}=27Cn^{1/4}$. By choice of $P$, we know that |$|\{x\in X:\chi(xy)=\text{blue}\}|<2C\sqrt{n}$ for every $y\in Y$, which implies that $|X_{1}|<2C\sqrt{n}|Y_{1}|<54C^{2}n^{3/4}$. Now, either $Y_{1}=\emptyset=X_{1}$, or, since $n$ is sufficiently large, we have
$\frac{|X_{0}|}{|Y_{1}|}=\frac{|X|-|X_{1}|}{|Y_{1}|}>\frac{n-(1+\alpha_{n})n^{1/2}-54C^{2}n^{3/4}}{27Cn^{1/4}}>n^{1/2}>\frac{108C^{2}n^{3/4}}{(1-2\alpha_{n})n^{1/2}}>\frac{2|X_{1}|}{|Y_{0}|}.$
@@ Acknowledgements
The second author is supported by the London Mathematical Society and the Heilbronn Institute for Mathematical Research through an Early Career Fellowship. The third author is supported by the Martingale Foundation.

### v2

@@ 2 Preliminaries
Let $C=20^{4}$ and $n>C^{10}$. Define $\alpha_{n}=11C/\sqrt[4]{n}$, and suppose $\chi\in\mathcal{X}_{n}$ is such that $f(n,\chi)>\sqrt{n}$. By Proposition 3.4, we have $f(m)<\sqrt{m}+C$ for every $m\in\mathbb{N}$. Applying Lemma 3.2 with $C_{1}=C$ and $C_{2}=0$, we see that without loss of generality, there exists a blue path $P$ such that the size of $Y=[n]\setminus V(P)$ is at most $(1+\alpha_{n})\sqrt{n}$ and with the additional property that every vertex in $Y$ has at most $2(C+1)\sqrt{n}$ blue neighbours in $X=V(P)$. By Lemma 3.3, if $|Y_{0}|\leq(1-2\alpha_{n})\sqrt{n}$ or $|Y|\leq\sqrt{n}-1$, then $f(n,\chi)\leq\sqrt{n}$, a contradiction. Let us therefore assume that $|Y_{0}|>(1-2\alpha_{n})\sqrt{n}$ and $|Y|>\sqrt{n}-1$.
We define $H$ to be the bipartite graph on vertex set $X\cup Y$ whose edges are the red edges between $X$ and $Y$ under $\chi$. Let $Y_{1}=Y\setminus Y_{0}$, $X_{0}=\{x\in X:d_{H}(x)=|Y|\}$, and $X_{1}=X\setminus X_{0}$. We have $|Y_{1}|=|Y|-|Y_{0}|<(1+\alpha_{n})\sqrt{n}-(1-2\alpha_{n})\sqrt{n}=33Cn^{1/4}$. By choice of $P$, we know that $|\{x\in X:\chi(xy)=\text{blue}\}|<2(C+1)\sqrt{n}$ for every $y\in Y$, which implies that $|X_{1}|<2(C+1)\sqrt{n}|Y_{1}|<67C^{2}n^{3/4}$. Now, either $Y_{1}=\emptyset=X_{1}$, or, since $n$ is sufficiently large, we have
$\frac{|X_{0}|}{|Y_{1}|}=\frac{|X|-|X_{1}|}{|Y_{1}|}>\frac{n-(1+\alpha_{n})n^{1/2}-67C^{2}n^{3/4}}{33Cn^{1/4}}>n^{1/2}>\frac{134C^{2}n^{3/4}}{(1-2\alpha_{n})n^{1/2}}>\frac{2|X_{1}|}{|Y_{0}|}.$
@@ Acknowledgements
The second author is supported by the London Mathematical Society and the Heilbronn Institute for Mathematical Research through an Early Career Fellowship. The third author is supported by the Martingale Foundation. We thank Zach Hunter and Frank Oberson for their comments on the paper.